OpenStudy (anonymous):

(2+3i)(2-3i) simplify. Plz help!!

OpenStudy (anonymous):

\[(2 + 3i)(2 - 3i) = 4 - 6i + 6i - 9i^{2} = 4 - 9i^{2} = 4 + 9 = 13\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You're welcome :)